# R-SCoRe: Revisiting Scene Coordinate Regression for Robust Large-Scale Visual Localization

### v1

@@ Abstract
Learning-based visual localization methods that use scene coordinate regression (SCR) offer the advantage of smaller map sizes. However, on datasets with complex illumination changes or image-level ambiguities, it remains a less robust alternative to feature matching methods. This work aims to close the gap. We introduce a covisibility graph-based global encoding learning and data augmentation strategy, along with a depth-adjusted reprojection loss to facilitate implicit triangulation. Additionally, we revisit the network architecture and local feature extraction module. Our method achieves state-of-the-art on challenging large-scale datasets without relying on network ensembles or 3D supervision. On Aachen Day-Night, we are 10$\times$ more accurate than previous SCR methods with similar map sizes and require at least 5$\times$ smaller map sizes than any other SCR method while still delivering superior accuracy. Code will be available at: https://github.com/cvg/scrstudio.
@@ 1 Introduction
Visual localization is the task of estimating the 6-DoF pose of a camera in a known scene with a query image. It is a fundamental problem in computer vision, with applications in augmented/virtual reality (AR/VR), autonomous driving, and robotics.
@@
SCR methods follow a structure-based paradigm similar to feature matching, i.e., estimating pose from 2D-3D correspondences, but replacing explicit matching with regressing the correspondences directly. They are usually limited to small scenes [9] and have yet to match feature matching methods in terms of accuracy and robustness. Recent advancements [74] extend SCR to large-scale scenes using a single model. However, its performance is still not on par with feature matching methods, especially in complex scenes with challenging illumination changes [60, 63].
In this work, we conduct a detailed analysis of the design principles behind the SCR framework, including local and global encoding, network architecture, and training strategies. Based on this analysis, we propose revisiting SCR to enhance the robustness and accuracy of SCR methods for large-scale visual localization tasks. As shown in Fig. 1, our robust SCR (R-SCoRe) improves the night-time localization accuracy for SCR methods to 64.3% under the (0.25m, 2°) threshold on the Aachen Day-Night dataset, all with a map size of only 47MB. R-SCoRe significantly outperforms previous SCR methods and achieves accuracy comparable to feature matching techniques.
Tab. 1 summarizes the practicability of R-SCoRe on complex large-scale scenes. While feature matching methods are also accurate, their map size can be prohibitively large, sometimes by more than two orders of magnitude [56, 57, 20]. Compared to SCR methods with similarly small map sizes [9, 74], R-SCoRe is at least one order of magnitude more accurate. While we still clearly outperform other SCR methods, we maintain fast inference and significantly smaller map sizes – all without the need for depth supervision.
Our contributions are summarized as follows:
  • We propose learning a global encoding and performing the data augmentation based on the covisibility graph. To address the ambiguity of image retrieval features in large-scale and complex scenes, we use multiple global hypotheses during testing.
  • To unbias the network from neglecting near points, we introduce a depth-adjusted reprojection loss and show that this allows for accurate localization without ground truth coordinate supervision.
@@ 2 Related Work
Feature Matching. Most state-of-the-art visual localization methods rely on feature matching [56, 57, 20, 62]. These methods typically adopt a structure-based localization paradigm, establishing 2D-3D correspondences between keypoints in a query image and 3D points in a scene. Camera pose is solved with geometric constraints, often utilizing a Perspective-n-Point (PnP) solver [27, 50] within a RANSAC framework [26, 2, 18] to effectively manage outliers. These methods commonly construct a Structure-from-Motion (SfM) map of the scene, which contains both 3D points and their corresponding descriptors [65, 25, 20, 54, 23]. To efficiently establish 2D-3D matches, these methods often implement a two-level hierarchical approach [56]. Initially, potentially relevant database images are retrieved using image retrieval techniques [1, 52, 84]. This is followed by 2D-2D matching with the query image [56, 57, 41].
@@
Recently, several approaches have been proposed to improve the scalability and performance of SCR for visual localization in large-scale scenes. These methods often rely on ground truth 3D coordinates and aim to handle large scenes by dividing them into smaller segments, such as spatial regions [5], voxels [70], or hierarchical clusters [38, 76]. Recent advancements have explored alternatives that do not require ground truth 3D supervision. For example, ACE [9] uses reprojection loss only for training. GLACE [74] further introduces a global encoding mechanism that eliminates the need for scene segmentation. Despite these improvements, these methods still encounter limitations under challenging conditions, such as significant changes in illumination.
@@ 3 Method
In this section, we first briefly review visual localization with Scene Coordinate Regression (SCR). Subsequently, we analyze each component and propose our R-SCoRe pipeline, shown in Fig. 2, to improve the performance and robustness of SCR.
@@ 3.1 Preliminaries
Visual Localization. Given a test image $I_{\text{test}}$ with known intrinsic $\mathtt{K}_{\text{test}}$, the objective is to estimate its extrinsic, which is the rigid transformation $[\mathtt{R}_{\text{test}}|\mathbf{t}_{\text{test}}]$ from world coordinate to camera coordinate. The scene is typically given by a set of training images $I_{\text{train}}$ with corresponding ground truth poses $[\mathtt{R}_{\text{train}}|\mathbf{t}_{\text{train}}]$ and intrinsics $\mathtt{K}_{\text{train}}$.
Scene Coordinate Regression. The SCR pipeline employs a neural network $f$ to directly regress the 3D coordinate $y$ for each 2D keypoint $x$ with feature $\mathbf{F}(x)$:
$y=f(\mathbf{F}(x)).$ (1)
Without the need to store large point clouds with descriptors, SCR methods implicitly represent the scene with a neural network, which usually results in a smaller map size.
Scalable SCR without 3D ground truth. Recent advances [9, 74] allow SCR to work on large-scale scenes without ground truth 3D supervision. To reduce ambiguities in large scenes, GLACE [74] combines a local patch-level encoding with an image-level global encoding as the keypoint feature $\mathbf{F}(x)$. The local encoder is pretrained as a shared DSAC* [6] backbone, following [9]. The global encoder leverages a pretrained image retrieval model [84] with a specific augmentation strategy to prevent overfitting to trivial solutions, cf. [74]. To accelerate training, all features are precomputed and stored in a buffer in GPU memory, from which a random sample is drawn for each batch.
Without ground truth scene coordinates, the output 3D point $y$ is reprojected to image space with the ground truth pose $\mathtt{R},t$ and intrinsics $\mathtt{K}$, and compared to the keypoint location $x$ in 2D:
@@
The reprojection error is fed into a dynamic robust loss function, cf. ACE [9], to focus on points that can be accurately regressed:
@@
where $\bar{y}$ is computed by the inverse projection of the pixel $x$ using the ground truth pose, intrinsics, and a fixed target depth $d_{\text{target}}$.
Figure 2: R-SCoRe pipeline. Adjustments to existing SCR pipelines, e.g. [74], are outlined in green. We learn global encodings aligned with the covisibility graph using Node2Vec [28] and a simple but effective data augmentation technique. We augment the pipeline with an additional refinement module and introduce a depth-adjusted reprojection loss for the coarse output $y_{0}$, mitigating bias toward distant points. Additionally, we explore alternative local encodings using off-the-shelf feature extractors [25, 69].
@@ 3.2 Network Architecture
We adopt the MLP architecture and position decoder from GLACE [74] and scale the network width according to scene complexity. As illustrated in Fig. 2, we also introduce a refinement module at the end of the network, which adjusts the final output by predicting an offset from the intermediate prediction $y_{0}$ to the final prediction $y$. The coarse coordinate $y_{0}$ is reintroduced into the refinement module through positional encoding using sine and cosine functions with periods ranging from 0.5 to 2048, which is then directly added to the intermediate feature. This empirically improves training stability and allows the network to achieve lower training reprojection errors more rapidly.
@@ 3.3 Input Encoding
Analysis. Previous studies [74, 48] suggest that SCR with reprojection loss supervision performs implicit triangulation where reprojection constraints are grouped by input similarity. Therefore, the desired properties of input encodings are as follows: positive pairs observing the same points should produce similar features, while negative pairs observing distinct points should yield clearly distinguishable features. Additionally, it is preferable for the encodings to be low-dimensional to minimize memory requirements.
@@
Figure 3: Comparison of global encodings. Aligning the learning of the global encodings to the topology of the covisibility graph (Node2Vec [28]) helps to disambiguate covisibile and non-covisible pairs, (a) right, and allows to predict covisibility by feature distance (b).
Covisibility Graph Based Global Encoding. Image-level global features should distinguish between covisible and non-covisible image pairs, i.e. whether the images are viewing the same part of the scene, to resolve ambiguities in local encodings. Although global encodings with image-level receptive fields can help, their low dimensionality may be insufficient to resolve ambiguities in complex environments, as shown in Fig. 3. This limitation can lead to imperfect grouping of reprojection constraints during training, thereby impairing the effectiveness of implicit triangulation. Furthermore, we point out that the learned SCR function may lack (Lipschitz) smoothness [34] w.r.t. the global encodings if adapted naively, e.g. minor variations in the global encoding can result in significant shifts in corresponding 3D points and consequently reduce generalization at test time.
To address these issues, we propose to directly learn embeddings aligned with the covisibility graph’s topology using Node2Vec [28], which samples sequences with weighted random walk and optimizes node embeddings with a Skip-gram [45] objective. For training images, the covisibility graph is easily available. It can be estimated from the frustum overlap of ground truth poses or through more accurate feature matching (see the supplementary for more details). At test time, however, covisibility information is unknown, so we propose generating multiple global encoding hypotheses by retrieving the nearest training images using NetVLAD [1] features. The global encoding of each retrieved image serves as a hypothesis, and we select the hypothesis with the maximum inliers for the final localization result.
@@ 3.4 Output Supervision
Selective Application of Depth Adjustment. The bias towards distant points may sometimes actually be beneficial, as underconstrained points can be pushed farther along the ray, making them easier to identify as outliers during test time. Therefore, we apply our depth-adjusted reprojection loss only to the intermediate coarse scene coordinate output $y_{0}$ during training (Fig. 4), and retain the original reprojection loss for the final output. To mitigate the concentration of the supervision signal on points with low projection error, we replace $\tanh$ with the Geman-McClure [3] robust loss function which has a heavier tail than $\tanh$:
@@
where $\bar{y}$ is the pseudo ground truth computed by the inverse projection of the pixel given the depth, pose and intrinsic.
@@ 4.1 Datasets
Hyundai Department Store. It is a large-scale indoor visual localization benchmark, covering three floors (B1, 1F, 4F) of a department store, with each floor consisting of multiple sequences captured over four months, spanning an area of approximately 10,000 $m^{2}$. The dataset presents challenges beyond its large scale, including dynamic objects, illumination changes, and textureless regions. B1 is particularly challenging as the training images are captured under low-lighting conditions, while the query images are brightly illuminated. The dataset includes 44,283 training images and 5,927 test images.
@@ 4.2 Benchmark Results
Aachen Day-Night. As shown in Tab. 2, R-SCoRe enhances the performance of scene coordinate regression (SCR) based methods [5, 74, 76, 38], achieving competitive results with a single low-map-size model without the need for ground truth depth supervision. While R-SCoRe is competitive with the best performing method, HLoc [56, 57, 20], we forfeit some ground at highest accuracy. However, R-SCoRe demands 170$\times$ less memory to store the map. This huge gap could already render SCR based methods as an attractive alternative for some applications. Most other feature based methods (FM) also deliver significantly larger maps. While delivering comparable performance for the Aachen Day dataset, they all fall behind R-SCoRe on the Aachen Night dataset. The only FM method [44] with a comparable map size is outperformed on all metrics, e.g. [44] is 4-5$\times$ worse in accuracy on the night dataset. Compared to other SCR methods that work without depth supervision [9, 74] R-SCoRe is 10$\times$ superior in accuracy. The so far most accuracte SCR based method [70] produces large maps (27$\times$ larger) and is prohibitively slow in inference. The next most accurate SCR method, [76] is outperformed by 46% at night and highest threshold, while R-SCoRe maintains a 6$\times$ smaller map size – without the need for depth supervision. We observe a small gain in performance if we utilize depth for supervision of R-SCoRe.
@@
Hyundai Department Store. As shown in Tab. 3, R-SCoRe again significantly outperforms the SCR based methods [5, 74], including recent ensemble networks [5, 9]. Furthermore, compared to the state-of-the-art feature matching based localization [56] we achieve competitive results with a single low-map-size model and again forfeit some ground at the highest accuracy threshold. However, our model is at least three orders of magnitude smaller for either feature [23, 54] incorporated into [56], which can be a valuable advantage in practice. Recall that ground truth depth supervision is not necessary for R-SCoRe, but if available, it can also enhance performance further. The B1 scene exhibits strong illumination changes and we observe significantly better performance when using local encodings from Dedode instead of LoFTR [69].

### v2

@@ Abstract
Learning-based visual localization methods that use scene coordinate regression (SCR) offer the advantage of smaller map sizes. However, on datasets with complex illumination changes or image-level ambiguities, it remains a less robust alternative to feature matching methods. This work aims to close the gap. We introduce a covisibility graph-based global encoding learning and data augmentation strategy, along with a depth-adjusted reprojection loss to facilitate implicit triangulation. Additionally, we revisit the network architecture and local feature extraction module. Our method achieves state-of-the-art on challenging large-scale datasets without relying on network ensembles or 3D supervision. On Aachen Day-Night, we are 10$\times$ more accurate than previous SCR methods with similar map sizes and require at least 5$\times$ smaller map sizes than any other SCR method while still delivering superior accuracy. Code is available at: https://github.com/cvg/scrstudio.
@@ 1 Introduction
Visual localization is the task of estimating the 6-DoF pose of a camera in a known scene with a query image. It is a fundamental problem in computer vision, with applications in augmented reality, autonomous driving, and robotics.
@@
SCR methods follow a structure-based paradigm similar to feature matching, i.e., estimating pose from 2D-3D correspondences but replacing explicit matching with regressing the correspondences directly. They are usually limited to small scenes [9] and have yet to match feature matching methods in terms of accuracy and robustness. Recent advances [75] extend SCR to large-scale scenes using a single model. However, its performance is still not on par with feature matching methods, especially in complex scenes with challenging illumination changes [60, 63].
In this work, we conduct a detailed analysis of the design principles behind the SCR framework, including local and global encoding, network architecture, and training strategies. Based on this analysis, we propose to revisit SCR to enhance the robustness and accuracy of SCR methods for large-scale visual localization tasks. As shown in Fig. 1, our robust SCR (R-SCoRe) improves the night-time localization accuracy for SCR methods to 64.3% under the (0.25m, 2°) threshold on the Aachen Day-Night dataset, all with a map size of only 47MB. R-SCoRe significantly outperforms previous SCR methods and achieves accuracy comparable to feature matching techniques.
Tab. 1 summarizes the practicability of R-SCoRe in complex large-scale scenes. While feature matching methods are also accurate, their map size can be prohibitively large, sometimes more than two orders of magnitude [56, 57, 20]. Compared to SCR methods with similarly small map sizes [9, 75], R-SCoRe is at least one order of magnitude more accurate. While we still clearly outperform other SCR methods, we maintain fast inference and significantly smaller map sizes – all without the need for scene-specific depth supervision. Our contributions are as follows.
  • We propose learning a global encoding and performing data augmentation based on the covisibility graph. To address the ambiguity of image retrieval features in complex large-scale scenes, we used multiple global hypotheses during testing.
  • To unbias the network from neglecting near points, we introduce a depth-adjusted reprojection loss and show that this allows for accurate localization without scene-specific ground truth coordinate supervision.
@@ 2 Related Work
Feature Matching. Most state-of-the-art visual localization methods rely on feature matching [56, 57, 20, 62]. These methods typically adopt a structure-based paradigm, establishing 2D-3D correspondences between keypoints in a query image and 3D points in a scene. Camera pose is solved with geometric constraints, often a Perspective-n-Point (PnP) solver [27, 50] within a RANSAC framework [26, 2, 18] to effectively manage outliers. These methods commonly construct a Structure-from-Motion (SfM) map of the scene, which contains both 3D points and their descriptors [65, 25, 20, 54, 23]. To efficiently establish 2D-3D matches, they often follow a two-level approach [56]. First, potentially relevant database images are retrieved using image retrieval techniques [1, 52, 85]. This is followed by 2D-2D matching with the query image [56, 57, 41].
@@
(a) Workflow Overview
(b) Detailed Pipeline
Figure 2: R-SCoRe pipeline. (a) Following the SCR workflow in [75], we concatenate patch-level local encodings with image-level global encodings as input to a scene-specific MLP. (b) We learn contrastive global encodings from the covisibility graph using Node2Vec [28]. During training, global encodings are sampled from neighboring nodes for data augmentation. During inference, we retrieve global encodings from the $k$ nearest training images via NetVLAD [1] as hypotheses and select the one yielding the most RANSAC inliers. We enhance the SCR MLP with a refinement module and introduce a depth-adjusted reprojection loss to reduce bias toward distant points.
@@
Recently, several approaches have been proposed to improve the scalability and performance of SCR in large-scale scenes. These methods often rely on ground truth 3D coordinates and aim to handle large scenes by dividing them into smaller segments, such as spatial regions [5], voxels [71], or hierarchical clusters [38, 77]. Recent advancements have explored alternatives that do not require ground truth 3D supervision. For example, ACE [9] uses reprojection loss only for training. GLACE [75] further introduces a global encoding mechanism that eliminates the need for scene segmentation. Despite these improvements, these methods still encounter limitations under challenging conditions, such as significant changes in illumination.
@@ 3 Method
Our Scene Coordinate Regression (SCR) workflow is depicted in Fig. 2(a). During training, we have access to a covisibility graph to learn global encodings which are concatenated to the local encodings. During inference, we retrieve global encoding hypotheses from the $k$ nearest training images and predict 2D-3D correspondences: we run PnP for each hypothesis and select the one yielding the most RANSAC inliers. Fig. 2(b) illustrates our detailed R-SCoRe pipeline, where we split SCR into coarse and refinement blocks and introduce covisibility-based global encoding and data augmentation techniques.
@@ 3.1 Preliminaries
Visual Localization. Given a test image $I_{\text{test}}$ with known intrinsic $\mathtt{K}_{\text{test}}$, the goal is to estimate its extrinsic, i.e., the rigid transformation $[\mathtt{R}_{\text{test}}|\mathbf{t}_{\text{test}}]$ from world coordinate to camera coordinate. The scene is typically given by a set of training images $I_{\text{train}}$ with known ground truth poses $[\mathtt{R}_{\text{train}}|\mathbf{t}_{\text{train}}]$ and intrinsics $\mathtt{K}_{\text{train}}$.
Scene Coordinate Regression. The SCR pipeline employs a neural network $f$ to directly regress the 3D coordinate $y=f(\mathbf{F}(x))$ for each 2D keypoint $x$ with feature $\mathbf{F}(x)$. Without the need to store large point clouds with descriptors, SCR methods implicitly represent the scene with a neural network, which usually results in a smaller map size.
Scalable SCR without 3D ground truth. Recent advances [9, 75] allow SCR to scale to large scenes without scene-specific 3D supervision. To reduce ambiguities in large scenes, GLACE [75] (Fig. 2(a)) concatenates a local patch-level encoding with an image-level global encoding as the keypoint feature $\mathbf{F}(x)$. The local encoder is a pretrained DSAC* [6] backbone, following [9]. The global encoder uses a pretrained image retrieval model [85] with Gaussian noise augmentation to prevent overfitting to trivial solutions, cf. [75]. To accelerate training, all features are precomputed and buffered in GPU memory, from which a random sample is drawn for each batch.
Without ground truth scene coordinates, the output 3D point $y$ is reprojected with the ground truth pose $\mathtt{R},t$ and intrinsics $\mathtt{K}$, and compared to the keypoint location $x$ in 2D:
@@
Instead of explicitly grouping corresponding observations into tracks, this underconstrained supervision is applied to each independent prediction. Prior works [75, 48] suggest that implicit triangulation can still occur as the network tends to produce similar outputs for similar inputs.
The reprojection error is fed into a dynamic robust loss, cf. ACE [9], to focus on points accurately regressed:
@@
where $\bar{y}$ is computed by the inverse projection of the pixel $x$ using a fixed target depth $d_{\text{target}}$.
@@ 3.2 Network Architecture
We adopt the MLP architecture and position decoder from GLACE [75], scaling the network width with scene size. As illustrated in Fig. 2(b), we also introduce a refinement module at the end of the network, which adjusts the final output $y$ by predicting an offset from the intermediate prediction $y_{0}$. The coarse coordinate $y_{0}$ is reintroduced into the refinement module through positional encoding [70, 46] using sine and cosine functions with periods ranging from 0.5 to 2048, which is added to the intermediate feature. This empirically improves training stability and allows the network to achieve lower training reprojection errors more rapidly.
@@ 3.3 Input Encoding
Analysis. In implicit triangulation, reprojection constraints are grouped based on input similarity. Therefore, the desired properties of input encodings are as follows: positive pairs observing the same points should produce similar features, while negative pairs observing distinct points should yield clearly distinguishable features. Additionally, it is preferable for the encodings to be low-dimensional to minimize memory requirements.
@@
Figure 3: Comparison of global encodings. Aligning the learning of global encodings with the covisibility graph topology (Node2Vec [28]) helps distinguish covisible and non-covisible pairs (a) and predict covisibility by feature distance (b).
Covisibility Graph Based Global Encoding. Image-level global features should distinguish between covisible and non-covisible image pairs, i.e. whether the images are viewing the same part of the scene, to resolve ambiguities in local encodings. Although global encodings with image-level receptive fields can help, they may still be insufficient to resolve ambiguities in complex environments, as shown in Fig. 3. This limitation can lead to imperfect grouping of reprojection constraints during training, thereby impairing the effectiveness of implicit triangulation. Furthermore, we point out that the learned SCR function may lack (Lipschitz) smoothness [34] w.r.t. the global encodings if adapted naively, e.g. minor variations in the global encoding can result in significant shifts in corresponding 3D points and consequently reduce generalization at test time.
To address these issues, we propose to directly learn embeddings aligned with the covisibility graph’s topology using Node2Vec [28], which samples sequences with weighted random walks and optimizes node embeddings with a Skip-gram [45] objective. For training images, the covisibility graph is easily available. It can be estimated from the frustum overlap of ground truth poses (see the supplementary for more details). At test time, however, covisibility information is unknown, so we propose generating multiple global encoding hypotheses by retrieving the nearest training images using NetVLAD [1] features. The global encoding of each retrieved image serves as a hypothesis, and we select the hypothesis yielding the maximum RANSAC inliers for the final localization result.
@@ 3.4 Output Supervision
Selective Application of Depth Adjustment. The bias towards distant points may sometimes actually be beneficial, as underconstrained points can be pushed farther along the ray, making them easier to identify as outliers during test time. Therefore, we apply our depth-adjusted reprojection loss only to the intermediate coarse scene coordinate output $y_{0}$ during training (Fig. 2(b)), and retain the original reprojection loss for the final output. To mitigate the concentration of the supervision signal on points with low projection error, we replace $\tanh$ with the Geman-McClure [3] robust loss function which has a heavier tail than $\tanh$:
@@
where $\bar{y}$ is the pseudo ground truth computed by the inverse projection of the pixel given the depth, pose, and intrinsic.
@@ 4.1 Datasets
Hyundai Department Store. It is a large-scale indoor visual localization benchmark, covering three floors of a department store. Each floor consists of multiple sequences captured over four months, spanning an area of approximately 10,000 $m^{2}$. It presents challenges beyond its large scale, including dynamic objects, illumination changes, and textureless regions. B1 is particularly challenging as the training images are captured under low-lighting conditions, while the query images are brightly illuminated. The dataset includes 44,283 training images and 5,927 test images.
@@ 4.2 Benchmark Results
Aachen Day-Night. As shown in Tab. 2, R-SCoRe enhances the performance of scene coordinate regression (SCR) based methods [5, 75, 77, 38], achieving competitive results with a single low-map-size model without the need for scene-specific depth supervision. While R-SCoRe is competitive with the best performing method, HLoc [56, 57, 20], we forfeit some ground at highest accuracy. However, R-SCoRe demands 170$\times$ less memory to store the map. This huge gap could already render SCR based methods as an attractive alternative for some applications. Most other feature based methods (FM) also deliver significantly larger maps. While delivering comparable performance for the Aachen Day dataset, they all fall behind R-SCoRe on the Aachen Night dataset. The only FM method [44] with a comparable map size is outperformed on all metrics, e.g. [44] is 4-5$\times$ worse in accuracy on the night dataset. Compared to other SCR methods that work without depth supervision [9, 75] R-SCoRe is 10$\times$ superior in accuracy. The so far most accuracte SCR based method [71] produces large maps (27$\times$ larger) and is prohibitively slow in inference. The next most accurate SCR method, [77] is outperformed by 46% at night and highest threshold, while R-SCoRe maintains a 6$\times$ smaller map size – without the need for depth supervision. We observe a small gain in performance if we utilize depth for supervision of R-SCoRe.
@@
Hyundai Department Store. R-SCoRe again significantly outperforms the SCR based methods [5, 75], including recent ensemble networks [5, 9], see Tab. 3. Compared to the state-of-the-art feature matching based localization [56] we achieve competitive results with a single low-map-size model and forfeit some ground at the highest accuracy threshold. However, our model is at least three orders of magnitude smaller for either feature [23, 54] incorporated into [56], which can be a valuable advantage in practice. Recall that depth supervision is not necessary for R-SCoRe, but if available, it can also enhance performance further. The B1 scene exhibits strong illumination changes and we observe significantly better performance when using local encodings from Dedode instead of LoFTR [69].